# DragPoser: Motion Reconstruction from Variable Sparse Tracking Signals via Latent Space Optimization

### v1

@@ 5.2. Dynamic Constraints Evaluation
5.3. Ablation Study
@@
Ablation Joint Error End-Effector Error
@@
No $\mathcal{L}_{C}$ 2.65(2.79) 8.39(9.52) 9.53(13.3) 1.50(1.13)
No $\mathcal{T}$ 3.05(4.39) 10.5(12.2) 13.1(19.0) 1.36(1.07)
Ours 2.18(2.70) 7.66(8.71) 10.7(15.1) 1.02(0.73)
Table 3. Results of the ablation study. The first row shows the accuracy metrics if removing the continuity loss $\mathcal{L}_{C}$ when training the Pose Autoencoder. The second row shows the impact of preventing the Pose Optimizer from using the output from the Temporal Predictor $\mathcal{T}$. As in previous tables, we report the joint errors and the end-effector errors.
@@
Our results indicate that the inclusion of $\mathcal{L}_{C}$ generally enhances all performance metrics by structuring the latent space more effectively. An exception is observed in the velocity metric; this is attributed to the fact that when the Continuity Loss is used, the optimization process needs fewer iterations to find poses that satisfy the constraints. Consequently, a lower learning rate might be necessary to avoid overshooting the target.
@@
Figure 6. Poses reconstructed using four sparse tracking signals (hip, head and hands) with (middle row) and without (last row) temporal predictor. Ground truth is colored in green. The Temporal Predictor guides the optimization process to produce temporally-coherent motion even in high ambiguity scenarios, thus producing a full walking cycle during motion (left). It also prevents implausible poses by guiding the optimization away from invalid latent space regions (right).
@@
The purpose of the Temporal Predictor is twofold. On the one hand, it ensures that predicted latent vectors remain within the manifold of valid poses. On the other hand, it learns the temporality within the latent space to guide the optimization process in high-ambiguity scenarios. In its absence, we observe a significant decline in all performance metrics, with Joint Error notably impacted due to its direct influence on the overall pose quality and not just the optimization of end-effector positioning.
@@
The effectiveness of the Temporal Predictor is even more pronounced in complex motions like push-ups. While the optimization process alone can satisfy end-effector constraints, the Temporal Predictor plays a crucial role in achieving temporally coherent poses. Finally, the last column demonstrates the Temporal Predictor’s ability to avoid invalid regions in the latent space: when disabled, the optimization process traverses invalid regions of the pose manifold, becomes trapped, and ultimately reconstructs an implausible pose.
6. Conclusions
In conclusion, our work addresses the critical challenge of achieving high-quality motion reconstruction with a reduced number of input devices, aiming to make motion capture more accessible to a wider audience. Our proposed method provides a perfect equilibrium between end-effector accuracy and pose quality. Through a pose optimization process within a structured latent space, our approach simultaneously achieves superior real-time end-effector position accuracy compared to state-of-the-art IK solutions, and higher pose quality compared to advanced learning approaches like AvatarPoser and SparsePoser. In addition, our system requires only one-time training on a large human motion dataset, allowing for dynamic constraint definition through losses during runtime. Unlike previous methods limited to specific tracker configurations, our approach can adapt to flexible tracker setups and can handle potential data loss due to occlusions or malfunctions. In such instances, while there may be a slight reduction in pose or end-effector accuracy for certain configurations, the structured latent space and temporal predictor enable our method to generate plausible and smooth animations. Overall, our approach not only expands accessibility to motion capture but also proves to be versatile and effective in handling real-world challenges, showcasing its potential for practical applications.
@@ 6. Conclusions
Despite the notable achievements of our motion reconstruction system, there are certain limitations that need further consideration. Firstly, in setups with only 3 sensors available, the lack of pelvis/root position tracking has a negative impact on the quality of our motion reconstruction. To overcome this limitation and improve the overall system’s performance, a promising approach would be to integrate a global position prediction module. Additionally, while our system is trained with various user dimensions and encodes poses using dual quaternions, the lack of comprehensive data specifically focusing on diverse user dimensions means that we do not explicitly validate its performance in such scenarios. Another potential limitation arises in instances where the pose optimization process may inadvertently move the latent vector outside the valid pose manifold. Although the inclusion of the Temporal Predictor effectively encodes valid regions, there remains a slim possibility of reconstructing invalid poses. While this occurrence is infrequent, it is essential to acknowledge and address this potential vulnerability. Furthermore, it is important to note that computing the gradient of the decoder can be time-consuming compared to standard forward prediction. Despite this computational demand, the optimization process typically operates with a few iterations, ensuring relatively fast performance.
@@
Future research directions may involve exploring the integration of various sensor types, such as computer vision, within the dynamic constraints of the pose optimizer. This extension would allow us to have additional sensor data, contributing to a more robust and versatile motion capture system. To further enhance the capabilities of our pose optimizer, we aim to introduce additional constraints, such as look-at functionality or end-effector pose preservation during editing. These improvements could provide more nuanced control over reconstructed motion, meeting diverse user needs and preferences. In the pursuit of refining the latent space structure, we plan to investigate alternative generative approaches. This includes exploring methods that prevent unbounded properties and ensure there are no empty spaces between valid regions. Such exploration could result in a more compact and well-organized latent space, thus enhancing the system’s efficiency and performance. Lastly, despite initially dismissing the idea of directly encoding temporality in the latent space due to increased complexity, we recognize it as an intriguing avenue for future research. Revisiting this concept and devising efficient strategies to manage the expanded complexity could potentially unlock new dimensions of temporal coherence and further improve the realism of motion reconstructions.
Acknowledgements.
This work has received funding from the European Union’s Horizon 2020 research and innovation programme under HORIZON-CL4-2022-HUMAN-01 grant agreement No 101093159 (XR4ED), and from MCIN/AEI/10.13039/501100011033/FEDER ”A way to make Europe”, UE (PID2021-122136OB-C21). This project has also received funding from the European Union’s Horizon 2020 Research and Innovation Programme under Grant Agreement No 739578 and the Government of the Republic of Cyprus through the Deputy Ministry of Research, Innovation and Digital Policy. Jose Luis Ponton was also funded by the Spanish Ministry of Universities (FPU21/01927).

### v2

@@ 3 DoF Position
5.3 Performance
[…]
To assess computational demands, we recorded the average time per frame/pose reconstruction compared to other methods. FinalIK, relying on optimized C# code, is the fastest at 0.14 ms. Among data-driven methods, AvatarPoser and SparsePoser lead at 2.15 ms and 15.6 ms, respectively, due to their purely forward-pass execution. HuMoR, hampered by sequence-level optimization due to its latent space design, requires 1567 ms per frame. In contrast, DragPoser, while incorporating optimization, achieves 25.3 ms per pose, making it suitable for interactive use. Note that we used a maximum iteration count of 40, which is almost never required due to early constraint fulfillment. DragPoser thus bridges the gap between real-time performance and the flexibility of latent space optimization, offering a cost-effective alternative to methods like HuMoR while approaching the speed of forward-pass approaches.
5.4 Ablation Study
@@
Table 3: Results of the ablation study. In the first row experiment, the continuity loss $\mathcal{L}_{C}$ is removed when training the Pose Autoencoder. In the second row experiment the Temporal Predictor $\mathcal{T}$ is not used. We report the joint errors and the end-effector errors.
Ablation End-Effector Error Joint Error
@@
No $\mathcal{L}_{C}$ 1.50(1.13) 8.39(9.52) 9.53(13.3) 2.65(2.79)
No $\mathcal{T}$ 1.36(1.07) 10.5(12.2) 13.1(19.0) 3.05(4.39)
Ours 1.02(0.73) 7.66(8.71) 10.7(15.1) 2.18(2.70)
@@
Our results indicate that including $\mathcal{L}_{C}$ generally enhances all performance metrics by structuring the latent space more effectively. An exception is observed in the velocity metric; this is attributed to the fact that when the Continuity Loss is used, the optimization process needs fewer iterations to find poses that satisfy the constraints. Consequently, a lower learning rate might be necessary to avoid overshooting the target.
@@
Figure 5: Poses reconstructed using four sparse tracking signals (hip, head and hands) with (middle row) and without (last row) temporal predictor. Ground truth is colored in green. The Temporal Predictor guides the optimization process to produce temporally-coherent motion even in high ambiguity scenarios, thus producing a full walking cycle during motion (left). It also prevents implausible poses by guiding the optimization away from invalid latent space regions (right).
@@
The purpose of the Temporal Predictor is twofold. On the one hand, it ensures that predicted latent vectors remain within the manifold of valid poses. On the other hand, it learns the temporality within the latent space to guide the optimization process in high-ambiguity scenarios. In its absence, we observe a significant decline in all metrics, with Joint Error notably impacted due to its direct influence on the overall pose quality and not just the optimization of end-effector positioning.
@@
The effectiveness of the Temporal Predictor is even more pronounced in complex motions like push-ups. While the optimization process alone can satisfy end-effector constraints, the Temporal Predictor plays a crucial role in achieving temporally coherent poses. Finally, the last column demonstrates the Temporal Predictor’s ability to avoid invalid regions in the latent space: when disabled, the optimization process traverses invalid regions of the pose manifold and ultimately reconstructs an implausible pose.
6 Conclusions
In conclusion, our work addresses the critical challenge of achieving high-quality motion reconstruction with a reduced number of input devices, aiming to make motion capture more accessible to a wider audience. Our proposed method provides a perfect equilibrium between end-effector accuracy and pose quality. Through a pose optimization process within a structured latent space, our approach simultaneously achieves superior real-time end-effector position accuracy compared to state-of-the-art IK solutions, and higher pose quality compared to advanced learning approaches like AvatarPoser, SparsePoser and HuMoR. In addition, our system requires only one-time training on a large human motion dataset, allowing for dynamic constraint definition through losses during runtime. Unlike previous methods limited to specific tracker configurations, our approach can adapt to flexible tracker setups and can handle potential data loss due to occlusions or malfunctions. In such instances, while there may be a slight reduction in pose or end-effector accuracy for certain configurations, the structured latent space and temporal predictor enable our method to generate plausible and smooth animations. Overall, our approach not only expands accessibility to motion capture but also proves to be versatile and effective in handling real-world challenges, showcasing its potential for practical applications.
@@ 6 Conclusions
Despite the notable achievements of our motion reconstruction system, certain limitations need further consideration. Firstly, in setups with only 3 sensors available, the lack of pelvis/root position tracking has a negative impact on the quality of our motion reconstruction. A promising approach would be to integrate a global position prediction module to overcome this limitation and improve the overall system’s performance. Additionally, while our system is trained with various user dimensions and encodes poses using dual quaternions, the lack of comprehensive data specifically focusing on diverse user dimensions means we do not explicitly validate its performance in such scenarios. Another potential limitation arises in instances where the pose optimization process may inadvertently move the latent vector outside the valid pose manifold. Although including the Temporal Predictor effectively encodes valid regions, the possibility of reconstructing invalid poses remains. This is primarily due to the weight used to balance between the Temporal Predictor and the optimization outputs. The lower the weight of the Temporal Predictor, the higher the probability of the optimization process of falling into invalid regions and getting stuck. Furthermore, it is important to note that computing the gradient of the decoder can be time-consuming compared to standard forward prediction. Despite this computational demand, the optimization process typically operates with a few iterations, ensuring relatively fast performance. Finally, DragPoser prioritizes high-frequency detail for accurate end-effector placement, which can sometimes introduce jitter. This can be mitigated by post-processing or by reducing the learning rate of the optimization process. Further research may explore alternative optimization strategies to refine this behavior, such as employing adaptive learning rate methods to control the optimization process better and reduce overshoot.
@@
Future research directions may involve exploring the integration of various sensor types, such as computer vision, within the dynamic constraints of the pose optimizer. This extension would allow us to have additional sensor data, contributing to a more robust and versatile motion capture system. To further enhance the capabilities of our pose optimizer, we aim to introduce additional constraints, such as look-at functionality or end-effector pose preservation during editing. These improvements could provide more nuanced control over reconstructed motion, meeting diverse user needs and preferences. To refine the latent space structure, we plan to investigate alternative generative approaches. This includes exploring methods that prevent unbounded properties and ensure no empty spaces between valid regions. Such exploration could result in a more compact and well-organized latent space, thus enhancing the system’s efficiency and performance. Lastly, despite initially dismissing the idea of directly encoding temporality in the latent vectors due to increased complexity, we recognize it as an intriguing avenue for future research. Currently, we encode this information in an external network. Revisiting this concept and researching efficient strategies to manage the expanded complexity could enhance the latent space representation.
7 Acknowledgements
This work has received funding from the European Union’s Horizon 2020 Research and Innovation Programme under HORIZON-CL4-2022-HUMAN-01 grant agreement No 101093159 (XR4ED) and grant agreement No 739578 (RISE), from MCIN/AEI/10.13039/501100011033/FEDER “A way to make Europe”, UE (PID2021-122136OB-C21), from the Department of Research and Universities of the Government of Catalonia (2021 SGR 01035), and the Government of the Republic of Cyprus through the Deputy Ministry of Research, Innovation and Digital Policy. Jose Luis Ponton was also funded by the Spanish Ministry of Universities (FPU21/01927), and Eduard Pujol by Universitat Politècnica de Catalunya and Banco Santander (FPI-UPC predoctoral grant).
[…]